Use the Laplace transform to solve the initial value problem: y + 4y = cos(2t), y(0) = 0, y(0 = 1.
Answer #1

3 y + gy = cos(2+) yo) = 0 1 y (0) = 1 Lly) + L (94) = 1/cos (2+)) s² Y(s) - Sy() -Y(o) + 9 Y(s) - S 52 +4 Y(s) s. + gy(5